OpenStudy (anonymous):

HELP!! PLEASE!! Will fan you if correct.. Louann and Carla received equal scores on a test made up of multiple choice questions and an essay. Louann got 14 multiple choice questions correct and received 24 points for her essay. Carla got 16 multiple choice questions correct and received 14 points for her essay. How many points was each multiple choice question worth? _________ points

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\text{let each multiple choice question worth= }x\\~\\ \text{score of louaann}\\~\\ =14x+24\\~\\ \text{score of carla}\\~\\ =16x+14\\~\\ \text{Louann and Carla received equal scores ,thus}\\~ \\ 14x+24=16x+14~~ }\end{align}\)

OpenStudy (mathmath333):

solve for x here

OpenStudy (anonymous):

So.. To get X I would add 14 and 24, and add 16 and 14? That kinda confused me a bit..

OpenStudy (mathmath333):

x is a variable u can only add variable with variable

OpenStudy (mathmath333):

for example \(x+8x\\=x(1+8)\\=9x\)

OpenStudy (anonymous):

So.. 16x+14x?

OpenStudy (mathmath333):

u cannot add like here \(1+7x=8\)---this is wrong

OpenStudy (mathmath333):

u have to shift \(14x\) to right hand side so it will become minus

OpenStudy (mathmath333):

\(14x+24=16x+14\\ 24-14=16x-14x\\ 10=x(16-14)\\ 10=2x\\ \dfrac{10}{2}=x\\ 5=x\)

OpenStudy (anonymous):

So it is x=5? I attempted it and got x=7 at first, i don't know how.. but that would be the answer to the question right?

OpenStudy (mathmath333):

show ur work

OpenStudy (anonymous):

Alright, one second.

OpenStudy (anonymous):

24 + 2x = 14 + 24 i didn't do the negatives.. I think?

OpenStudy (anonymous):

Could you check a few of my other answers? I think that they're right, not sure though.

OpenStudy (mathmath333):

look at the equation i made it is \(14x+24=16x+14\) compare it with yours

OpenStudy (anonymous):

Yeah, I see sorta. Pretty much, i put the wrong numbers?

OpenStudy (mathmath333):

yes lol

OpenStudy (anonymous):

Alright haha. Thanks, could you check some of my other answers? Not sure if they're correct I think they Might be..

OpenStudy (mathmath333):

yea

OpenStudy (anonymous):

Sorry, my computer died, i'll send the pictures here in a second.

OpenStudy (anonymous):

Are these correct?

OpenStudy (anonymous):

@mathmath333